OpenStudy (anonymous):

5x^6 * (-3x^3) help I will medal

OpenStudy (anonymous):

5x6(−3x3) $\mathrm{Remove\:parentheses}:\quad\left(-a\right)=-a$Remove parentheses: (−a)=−a $=-5x^6\cdot\:3x^3$=−5x6· 3x3 $\mathrm{Multiply\:the\:numbers:}\:5\cdot\:3=15$Multiply the numbers: 5· 3=15 $=-15x^6x^3$=−15x6x3 $\mathrm{Apply\:exponent\:rule}:\quad\:a^b\cdot\:a^c=a^{b+c}$Apply exponent rule: ab· ac=ab+c $x^6x^3=\:x^{6+3}=\:x^9$x6x3= x6+3= x9 $=-15x^9$=−15x9

OpenStudy (anonymous):

 Is this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just anwsered this

OpenStudy (anonymous):

$5x^6\left(-3x^3\right)$5x6(−3x3) $\mathrm{Remove\:parentheses}:\quad\left(-a\right)=-a$Remove parentheses: (−a)=−a $=-5x^6\cdot\:3x^3$=−5x6· 3x3 $\mathrm{Multiply\:the\:numbers:}\:5\cdot\:3=15$Multiply the numbers: 5· 3=15 $=-15x^6x^3$=−15x6x3 $\mathrm{Apply\:exponent\:rule}:\quad\:a^b\cdot\:a^c=a^{b+c}$Apply exponent rule: ab· ac=ab+c $x^6x^3=\:x^{6+3}=\:x^9$x6x3= x6+3= x9 $=-15x^9$=−15x9

OpenStudy (anonymous):

Yes I know @usercode3rror but I cannot understand what you said.

OpenStudy (anonymous):

=−​15x9

OpenStudy (anonymous):

5x6(−3x3) =−5x6· 3x3 −15x6x3\ $=-15x^9$=−​15x9

OpenStudy (anonymous):

HERE I will explain it in simple words. Multiply the "normal sized letters" together. 5 * -3 and add the exponents. 5 * -3 is what? 6 + 3 is what?

OpenStudy (anonymous):

britt i gave the anwser

OpenStudy (anonymous):

good job @usercode3rror

OpenStudy (anonymous):

;]

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

For those of you who are confused at what @usercode3rror said, he basically said multiply the numbers, and add the exponents. 3 * -5 = -15 6 + 3 (these are the exponents) = 9 so... -15x^9

OpenStudy (anonymous):

Now that I look at what you said @usercode3rror I can understand it...my computer didn't load everything you had said lol

OpenStudy (anonymous):

lol